# Theoretical Detailed Analyses for DC readout and a Fabri-Pérot gravitational-wave detector

### v1

@@ III General arguments for the DC-readout scheme
Here, we ignore the remaining term $O\left(\left|{{\mathfrak{B}}}\right|^{0}\right)$. The time-domain version of this signal operator is given by
@@
$=$ $\left\{\begin{array}[]{ccccc}1&&\mbox{for}&&a=0,\\ 0&&\mbox{for}&&a\neq 0,\end{array}\right$. (36)
@@
From the input-output relation which will be derived in Sec. VI, we can confirm the following expectation values:
@@
IV Input-Output relation of the Fabri-Pérot interferometer
In this section, we derive the input-output relation of the Fabri-Pérot interferometer, which is used in the setup of gravitational-wave detectors. The Fabri-Pérot interferometer consists of the beam splitter (BS), intermediate masses (ITMs), and the end mirrors (EMs). The injected laser to the Fabri-Pérot interferometer is separated into the directions of the $x$- and the $y$-arms. The laser is amplified between the ITM and EM. In gravitational-wave detectors, BS, ITMs, and EMs are suspended through the vibration-isolation system to measure the relative positions of these mirrors precisely. In these relative positions, gravitational-wave signals are included.
@@ IV.1 Mirror Displacements
To describe the relative positions of BS, ITMs, and EMs, we introduce the proper reference frame [25] whose origin is BS. As depicted in Fig. 1, we denote the coordinate values in the proper reference frame of ITM (XITM) and EM (XEM) along $x$-arm by $l_{x}+\hat{X}_{XITM}$ and $l_{x}+L+\hat{X}_{XEM}$, respectively. Here, $l_{x}$ is the distance between BS and XITM in a situation where there are no external forces, including the radiation pressure forces from the laser and gravitational-wave signals. $L$ is the distance between XITM and XEM in a situation where there are no external forces, including the radiation pressure forces from the laser and gravitational-wave signals. We regard $l_{x}$ and $L$ as classical distances. In addition to the $l_{x}$ and $L$, we introduce the quantum displacement $\hat{X}_{XITM}$ and $\hat{X}_{EM}$, which are induced by the injected laser and other forces, including the gravitational-wave signal. In this paper, we regard $\hat{X}_{XITM}$ and $\hat{X}_{EM}$ as quantum operators that describe the quantum mirror positions. In Sec. V, the operators $\hat{X}_{XITM}$ and $\hat{X}_{EM}$ are determined as the solution to the Heisenberg equation of motion.
@@
Based on the above setup, we consider the electric field of the laser through quantum electrodynamics.
@@ IV.3.1 Propagation between BS and ITMs
as the output electric field operators $\hat{E}_{c^{\prime}_{x}}(t)$ and $\hat{E}_{c^{\prime}_{y}}(t)$ from each arm. Here, we defined $l_{x}=:c\tau_{x}^{\prime}$ and $l_{y}=:c\tau_{y}^{\prime}$. In the Fourier space, we Eq. (62) is given by
@@ IV.3.2 Junction at ITMs and cavity propagation
Next, we consider the arm propagation of the laser along each arm. We denote the notation of the electric field of the electric field as depicted in Fig. 2.
Figure 2: Arm propagation in the Fabry-Pérot interferometer and photoelectric quadratures notations. The quadratures $\hat{f}$ and $\hat{f}^{\prime}$ are those of the laser incident from the BS to ITM and the laser from ITM to BS, respectively. The $\hat{g}$ and $\hat{g}^{\prime}$ are the quadratures for the lasers from ITM to the Fabri-Pérot cavity and from the Fabri-Pérot cavity to ITM, respectively. The quadratures $\hat{j}$ and $\hat{j}^{\prime}$ are quadratures for lasers that reach the EM and are reflected by the EM, respectively.
@@
If we have to consider the loss model of the imperfection of EM as in Ref. [8], we have to change Eq. (77).
Furthermore, the derivations of the electric field relations between $\hat{E}_{g_{x,y}}$ , $\hat{E}_{j_{x,y}}$, $\hat{E}_{j^{\prime}_{x,y}}$, and $\hat{E}_{g^{\prime}_{x,y}}$ are delicate. However, the physical effects are just free propagation of the laser. This is because the time-dependence of the relative positions $\hat{X}_{x}(t):=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}(t):=\hat{X}_{YEM}-\hat{X}_{YITM}$. We have to be careful to treat this time dependence. However, we apply the perfect reflection of EMs and the relation of $\hat{E}_{g_{x,y}}$ and $\hat{E}_{g^{\prime}_{x,y}}$ should be identical to the Michelson interferometer from the setup depicted in Fig. 2. Then, we obtain
@@
where $\tau=L/c$. Furthermore, through Eqs. (74), (75), (78), and (79), we obtain the relations
@@
To obtain the expressions of $\hat{E}_{g_{x,y}}$ in terms of $\hat{E}_{f_{x,y}}$, the treatment of the Fourier transformation is convenient. First, we consider the Fourier transformation of Eq. (80). Introducing the Fourier transformation of the displacement variable $\hat{X}_{x,y}(t)$ as
@@
we only considered the effects of the linear level of $\hat{X}_{x}$ in Eq. (80). Furthermore, through the inverse Fourier transformation, we obtain
@@ IV.4 Input-output relation with mirrors’ motion
The last integral in Eq. (94) which includes $\hat{Z}_{YITM}$ and $\hat{Z}_{XITM}$ are the retarded effects during the propagation of the laser from the intermediate masses to the beam splitter due to the modification of the proper distance between the beam splitter and the intermediate mirrors $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively. The third integral in Eq. (94) which includes $\hat{X}_{x}:=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}:=\hat{X}_{XEM}-\hat{X}_{XITM}$, is also the retarded effect during the propagation of the laser in the Fabri-Pérot cavities. The second integral in Eq. (94) are the retarded effect during the propagation from the beam splitter to the intermediate masses due to the modification of the proper distance between the beam splitter and the intermediate mirrors $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively. The first line is the direct shot noise of this interferometer. The last line in Eq. (94) symbolically represents the terms of the order of $O(\hat{X}_{x}^{2}$, $\hat{X}_{y}^{2}$, $\hat{X}_{XITM}^{2}$, $\hat{X}_{YITM}^{2})$.
@@
where $\hat{Z}_{com}(\omega)$ represents the $x$-arm and $y$-arm common motion of the relative motion between the end-mirror and the intermediate mass and $\hat{Z}_{dif}(\omega)$ represents the $x$-arm and $y$-arm differential motion of the relative motion the end-mirror and intermediate mass. Furthermore, we also define $\hat{Z}_{comITM,diffITM}(\omega)$ by
@@ IV.5 Coherent state of the optical fields
Here, Eq. (99) implies that the output operator $\hat{B}$ is given by the operators $\hat{A}$, $\hat{D}$, $\hat{Z}_{diff}$, $\hat{Z}_{com}$, $\hat{Z}_{diffITM}$, and $\hat{Z}_{comITM}$. Later, we see that the displacements $\hat{Z}_{diff}$, $\hat{Z}_{com}$, $\hat{Z}_{diffITM}$, and $\hat{Z}_{comITM}$ are given by $\hat{A}$ and $\hat{D}$ together with the gravitational-wave signal through the equations of motions for end-mirrors and the intermediate mirrors. Therefore, to discuss the information from the output operator $\hat{B}$, we have to specify the quantum states associated with the operators $\hat{A}$ and $\hat{D}$. The state associated with the operator $\hat{A}$ is the state of the electric field that is injected from the anti-symmetric port. At this anti-symmetric port, the photo-detector is located as in Fig. 1. On the other hand, the state associated with the operator $\hat{D}$ is the state of the electric field that is injected from the symmetric port. At this symmetric port, the light source exists as depicted in Fig. 1. The total state of photon in the output port $\hat{B}$, i.e., $\hat{b}$, is determined by the specification of the states associated with the operators $\hat{D}$ and $\hat{A}$, i.e., the annihilation and creation operators $(\hat{d},\hat{d}^{\dagger})$ and $(\hat{a},\hat{a}^{\dagger})$, respectively.
@@
$|\mbox{in}\rangle$ $=$ $\prod_{\omega}|\alpha(\omega)\rangle_{d}\otimes|0\rangle_{a}=\prod_{\omega}D_{d}[\alpha(\omega)]|0\rangle_{d}\otimes|0\rangle_{a}$ (100)
$=:$ $D_{d}|0\rangle_{d}\otimes|0\rangle_{a},$
$D_{d}$ $:=$ $\prod_{\omega}D_{d}(\alpha(\omega))$ (101)
@@
$D_{d}^{\dagger}\hat{d}(\omega)D_{d}=\hat{d}(\omega)+\alpha(\omega),$ (102)
$D_{d}^{\dagger}\hat{d}^{\dagger}(\omega)D_{d}=\hat{d}^{\dagger}(\omega)+\alpha^{*}(\omega),$ (103)
through the displacement operator $D_{d}$ defined by Eq. (101). Then, the operator $\hat{D}(\omega)$ defined by Eq. (61) is transformed as
$D_{d}^{\dagger}\hat{D}(\omega)D_{d}=\hat{D}_{c}(\omega)+\hat{D}_{v}(\omega),$ (104)
@@
To evaluate the output signal expectation value and its fluctuations from the input-output relation (99) under the coherent state of the quadrature $\hat{D}(\omega)$, the expression of $D_{d}^{\dagger}\hat{B}(\omega)D_{d}$ is useful instead of $\hat{B}(\omega)$, because we treat operators in the Heisenberg picture. Furthermore, we regard that $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$ are small correction due to the radiation-pressure noise and gravitational-wave signals. Since $\hat{D}_{v}(\omega)$ and $\hat{A}(\omega)$ also describe small fluctuations, we neglect the quadratic terms $\hat{D}_{v}(\omega)D_{d}^{\dagger}\hat{Z}_{*}(\Omega)D_{d}$ and $\hat{A}(\omega)D_{d}^{\dagger}\hat{Z}_{*}(\Omega)D_{d}$ in the expression of $D_{d}^{\dagger}\hat{B}(\omega)D_{d}$.
Operating $D_{d}^{\dagger}$ and $D_{d}$ to Eq. (99), substituting Eqs. (104) into this $D_{d}^{\dagger}$-$D_{d}$-operated version of Eq. (99), and applying the above approximation, then, we obtain the input-output relation as
$\!\!\!\!\!\!D_{d}^{\dagger}\hat{B}(\omega)D_{d}$ (107)
@@
$\quad\quad\times\left[i\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1})D_{d}-\cos(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1})D_{d}\right]$
@@
$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{com}(\omega-\omega_{1})D_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diff}(\omega-\omega_{1})D_{d}\right]$
@@
$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1})D_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1})D_{d}\right]$
@@
The input-output relation (107) is the most general input-output relation within our consideration. The first term in the first line of Eq. (107) is the leakage of the classical carrier field due to the phase offset $\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})$. The terms in the second line are the vacuum fluctuations which correspond to the shot noise in the conventional input-output relations in Refs. [8]. The terms in the third- and fourth-lines are the response of the intermediate mirror motion which includes the radiation pressure noise through the motions of the intermediate mirrors $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$. The terms in the fifth-, sixth-lines are the response of the relative motion of the end mirrors and the intermediate mirrors $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$. Finally, the term in the seventh- and eighth-lines are response of the intermediate mirror motion which includes the radiation pressure noise through the motion of the intermediate mirrors $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$. These come from the retarded effects which was explained after Eq. (94). The last line in Eq. (107) symbolically represents the terms of the order of $O(\hat{X}_{x}^{2}$, $\hat{X}_{y}^{2}$, $\hat{X}_{XITM}^{2}$, $\hat{X}_{YITM}^{2}$, $\hat{D}_{v}\hat{X}_{x}$, $\hat{D}_{v}\hat{X}_{y}$, $\hat{D}_{v}\hat{X}_{XITM}$, $\hat{D}_{v}\hat{X}_{YITM}$, $\hat{A}\hat{X}_{x}$, $\hat{A}\hat{X}_{y}$, $\hat{A}\hat{X}_{XITM}$, $\hat{A}\hat{X}_{YITM})$. The input-output relation (107) is one of the main results of this paper.
To conduct the actual evaluation of the input-output relation (107), we have to evaluate $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$ in some way. In the case of gravitational-wave detectors, $D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$ are evaluated through the equations of motions for the end-mirrors as discussed in the next section.

### v2

@@ III General arguments for the DC-readout scheme
Since we consider the situation where $|{{\mathfrak{B}}}|$ is so large that we can distinguish the effects due to the different order of $|{{\mathfrak{B}}}|$, we ignore the remaining term $O\left(\left|{{\mathfrak{B}}}\right|^{0}\right)$ because we want to measure the $O(|{{\mathfrak{B}}}|^{1})$ effect in the expectation value (26). The time-domain version of this signal operator is given by
@@
$=$ $\left\{\begin{array}[]{ccccc}1&\quad&\mbox{for}&\quad&a=0,\\ 0&\quad&\mbox{for}&\quad&a\neq 0,\end{array}\right$. (36)
@@
Although the final input-output relation will be derived in Sec. VI, we compare the final input-output relation (189) and Eqs. (19) and (20) above. Then, we obtain the expression of the operator $\hat{b}_{n}(\omega)$, more precisely ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$. Through this obtained expression of $\hat{b}_{n}(\omega)$, we can confirm the following expectation values:
@@
IV Input-Output relation of the Fabry-Pérot interferometer
In this section, we derive the input-output relation of the Fabry-Pérot interferometer, which is used in the setup of gravitational-wave detectors. The Fabry-Pérot interferometer consists of the Beam Splitter (BS), Input Test Masses (ITMs), and the End test Masses (EMs). The injected laser to the Fabry-Pérot interferometer is separated into the directions of the $x$- and the $y$-arms. The laser is amplified between the ITM and EM. In gravitational-wave detectors, BS, ITMs, and EMs are suspended through the vibration-isolation system to measure the relative positions of these mirrors precisely. In these relative positions, gravitational-wave signals are included through the equation of mirrors’ motion as extensively discussed in Sec. V.
@@ IV.1 Mirror Displacements
To describe the relative positions of BS, ITMs, and EMs, we introduce the proper reference frame [30] whose origin is BS. As depicted in Fig. 1, we denote the coordinate values in the proper reference frame of ITM (XITM) and EM (XEM) along $x$-arm by $l_{x}+\hat{X}_{XITM}$ and $l_{x}+L+\hat{X}_{XEM}$, respectively. Here, $l_{x}$ is the distance between BS and XITM, and $L$ is the distance between XITM and XEM in a situation where there are no external forces, including the radiation pressure forces from the laser and gravitational-wave signals. We regard $l_{x}$ and $L$ as classical distances. In addition to the $l_{x}$ and $L$, we introduce the quantum displacement $\hat{X}_{XITM}$ and $\hat{X}_{XEM}$, which are induced by the injected laser and other forces, including the gravitational-wave signal. In this paper, we regard $\hat{X}_{XITM}$ and $\hat{X}_{XEM}$ as quantum operators that describe the quantum mirror positions. In Sec. V, the operators $\hat{X}_{XITM}$ and $\hat{X}_{XEM}$ are determined as the solution to the Heisenberg equation of motion.
@@
Based on the above setup, we calculate the laser’s electric field using quantum electrodynamics.
@@ IV.3.1 Propagation between BS and ITMs
as the output electric field operators $\hat{E}_{c^{\prime}_{x}}(t)$ and $\hat{E}_{c^{\prime}_{y}}(t)$ from each arm. Here, we defined $l_{x}=:c\tau_{x}^{\prime}$ and $l_{y}=:c\tau_{y}^{\prime}$. In the Fourier space, Eq. (62) is given by
@@ IV.3.2 Junction at ITMs and cavity propagation
Next, we consider the laser’s propagation along each arm. We denote the notation of the electric field of the electric field as depicted in Fig. 2.
Figure 2: Arm propagation in the Fabry-Pérot interferometer and photoelectric quadratures notations. The quadratures $\hat{f}$ and $\hat{f}^{\prime}$ are those of the laser incident from the BS to ITM and the laser from ITM to BS, respectively. The $\hat{g}$ and $\hat{g}^{\prime}$ are the quadratures for the lasers from ITM to the Fabry-Pérot cavity and from the Fabry-Pérot cavity to ITM, respectively. The quadratures $\hat{j}$ and $\hat{j}^{\prime}$ are quadratures for lasers that reach the EM and are reflected by the EM, respectively.
@@
If we have to consider the loss model of the imperfection of EM as in Ref. [10], we have to change Eq. (77) as described in Ref. [10].
Furthermore, the derivations of the electric field relations between $\hat{E}_{g_{x,y}}$ , $\hat{E}_{j_{x,y}}$, $\hat{E}_{j^{\prime}_{x,y}}$, and $\hat{E}_{g^{\prime}_{x,y}}$ are delicate. However, the physical effects are just the free propagation of the laser. This is because the time-dependence of the relative positions $\hat{X}_{x}(t):=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}(t):=\hat{X}_{YEM}-\hat{X}_{YITM}$. We have to be careful to treat this time dependence correctly to estimate the retarded effects on the electric fields in the interferometer. Since we apply the perfect reflection of EMs, the relation of $\hat{E}_{g_{x,y}}$ and $\hat{E}_{g^{\prime}_{x,y}}$ is identical to the case of the Michelson interferometer from the setup depicted in Fig. 2. Then, we obtain
@@
where $\tau=L/c$. Moreover, through Eqs. (74), (75), (78), and (79), we obtain the relations
@@
To obtain the expressions of $\hat{E}_{g_{x,y}}$ in terms of $\hat{E}_{f_{x,y}}$, the treatment of the Fourier transformation is convenient, since we only consider the linear perturbation level with respect to $\hat{X}_{x,y}$ within this paper. First, we consider the Fourier transformation of Eq. (80). Introducing the Fourier transformation of the displacement variable $\hat{X}_{x,y}(t)$ as
@@
we only consider the effects of the linear level of $\hat{X}_{x}$ in Eq. (80), since we regard that the higher-order of $\hat{X}_{x}$ are negligible. Furthermore, through the inverse Fourier transformation, we obtain
@@ IV.4 Input-output relation with mirrors’ motion
The first line is the direct shot noise of this interferometer. The integral in the second and third lines in Eq. (94) is the retarded effect during the propagation from the beam splitter to the input test masses due to the modification of the proper distance between the beam splitter and the input test masses $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively. We note that these second and third lines do not include the propagation of the Fabry-Perot cavities between the ITMs and EMs, respectively. The integral in the fourth and fifth line in Eq. (94), which includes $\hat{X}_{x}:=\hat{X}_{XEM}-\hat{X}_{XITM}$ and $\hat{X}_{y}:=\hat{X}_{XEM}-\hat{X}_{XITM}$, is also the retarded effect during the propagation of the laser in the Fabry-Pérot cavities. The integral in the sixth and seventh line of Eq. (94), which includes $\hat{Z}_{YITM}$ and $\hat{Z}_{XITM}$, are the retarded effects during the propagation of the laser from the input test masses to the beam splitter due to the modification of the proper distance between the beam splitter and the input test masses $l_{x}+\hat{X}_{XITM}$ ( $l_{y}+\hat{X}_{YITM}$ ), respectively. The difference of these terms from the second and third lines in Eq. (94) is the fact that the sixth and seventh lines of Eq. (94) include the information of the propagation in the Fabry-Pérot cavities. The last line in Eq. (94) symbolically represents the terms of the order of $O(\hat{X}_{x}^{2}$, $\hat{X}_{y}^{2}$, $\hat{X}_{XITM}^{2}$, $\hat{X}_{YITM}^{2})$.
@@
where $\hat{Z}_{com}(\omega)$ represents the $x$-arm and $y$-arm common motion of the relative motion between the end test mass and the input test mass and $\hat{Z}_{diff}(\omega)$ represents the $x$-arm and $y$-arm differential motion of the relative motion the end test mass and input test mass. Furthermore, we also define $\hat{Z}_{comITM,diffITM}(\omega)$ by
@@ IV.5 Coherent state of the optical fields
Here, Eq. (99) implies that the output operator $\hat{B}$ is given by the operators $\hat{A}$, $\hat{D}$, $\hat{Z}_{diff}$, $\hat{Z}_{com}$, $\hat{Z}_{diffITM}$, and $\hat{Z}_{comITM}$. Later, we see that the displacements $\hat{Z}_{diff}$, $\hat{Z}_{com}$, $\hat{Z}_{diffITM}$, and $\hat{Z}_{comITM}$ are given by $\hat{A}$ and $\hat{D}$ together with the gravitational-wave signal through the equations of motions for end test masses and the input test masses. Therefore, to discuss the information from the output operator $\hat{B}$, we have to specify the quantum states associated with the operators $\hat{A}$ and $\hat{D}$. The state associated with the operator $\hat{A}$ is the state of the electric field that is injected from the anti-symmetric port. At this anti-symmetric port, the photodetector is located as in Fig. 1. On the other hand, the state associated with the operator $\hat{D}$ is the state of the electric field that is injected from the symmetric port. At this symmetric port, the light source exists as depicted in Fig. 1. The total state of photon in the output port $\hat{B}$, i.e., $\hat{b}$, is determined by the specification of the states associated with the operators $\hat{D}$ and $\hat{A}$, i.e., the annihilation and creation operators $(\hat{d},\hat{d}^{\dagger})$ and $(\hat{a},\hat{a}^{\dagger})$, respectively.
@@
$|\mbox{in}\rangle$ $=$ $\prod_{\omega}|\alpha(\omega)\rangle_{d}\otimes|0\rangle_{a}=\prod_{\omega}{{\mathfrak{D}}}_{d}[\alpha(\omega)]|0\rangle_{d}\otimes|0\rangle_{a}$ (100)
$=:$ ${{\mathfrak{D}}}_{d}|0\rangle_{d}\otimes|0\rangle_{a},$
${{\mathfrak{D}}}_{d}$ $:=$ $\prod_{\omega}{{\mathfrak{D}}}_{d}(\alpha(\omega))$ (101)
@@
These states are the reasonable realization of electric field operators that represent the injection of the coherent state laser from the light source, and no input sources except for vacuum fluctuations from the output port. We also note that the unitarity of the operation of the beam splitter on the electric field is guaranteed by the injection of the vacuum state associated with the operator $\hat{a}$ from the output port. Therefore, these setups are essential to the interferometer’s arguments.
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{d}(\omega){{\mathfrak{D}}}_{d}=\hat{d}(\omega)+\alpha(\omega),$ (102)
${{\mathfrak{D}}}_{d}^{\dagger}\hat{d}^{\dagger}(\omega){{\mathfrak{D}}}_{d}=\hat{d}^{\dagger}(\omega)+\alpha^{*}(\omega),$ (103)
through the displacement operator ${{\mathfrak{D}}}_{d}$ defined by Eq. (101). Then, the operator $\hat{D}(\omega)$ defined by Eq. (61) is transformed as
${{\mathfrak{D}}}_{d}^{\dagger}\hat{D}(\omega){{\mathfrak{D}}}_{d}=\hat{D}_{c}(\omega)+\hat{D}_{v}(\omega),$ (104)
@@
To evaluate the output signal expectation value and its fluctuations from the input-output relation (99) under the coherent state of the quadrature $\hat{D}(\omega)$, the expression of ${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega){{\mathfrak{D}}}_{d}$ is useful instead of $\hat{B}(\omega)$, because we treat operators in the Heisenberg picture. Furthermore, we regard that ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$ are small correction due to the radiation-pressure noise and gravitational-wave signals. Since $\hat{D}_{v}(\omega)$ and $\hat{A}(\omega)$ also describe small fluctuations, we neglect the quadratic terms $\hat{D}_{v}(\omega){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{*}(\Omega){{\mathfrak{D}}}_{d}$ and $\hat{A}(\omega){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{*}(\Omega){{\mathfrak{D}}}_{d}$ in the expression of ${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega){{\mathfrak{D}}}_{d}$.
Operating ${{\mathfrak{D}}}_{d}^{\dagger}$ and ${{\mathfrak{D}}}_{d}$ to Eq. (99), substituting Eqs. (104) into this ${{\mathfrak{D}}}_{d}^{\dagger}$-${{\mathfrak{D}}}_{d}$-operated version of Eq. (99), and applying the above approximation, then, we obtain the input-output relation as
$\!\!\!\!\!\!{{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega){{\mathfrak{D}}}_{d}$ (107)
@@
$\quad\quad\times\left[i\sin(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}-\cos(\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}\right]$
@@
$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}\right]$
@@
$\quad\quad\times\left[i\sin\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}-\cos\left((\omega+\omega_{1})\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\omega-\omega_{1}){{\mathfrak{D}}}_{d}\right]$
@@
The input-output relation (107) is the most general input-output relation within our consideration. The first term in the first line of Eq. (107) is the leakage of the classical carrier field due to the phase offset $\omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})$. The terms in the second line are the vacuum fluctuations that correspond to the shot noise in the conventional input-output relations in Refs. [10]. The terms in the third and fourth lines are the response of the input-test-mass motion, which includes the radiation pressure noise through the motions of the input test masses ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$. The terms in the fifth and sixth lines are the response of the relative motion of the end test masses and the input test masses ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$. Finally, the terms in the seventh and eighth lines are responses of the input-test-mass motion, which includes the radiation pressure noise through the motion of the input test masses ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$. These come from the retarded effects which were explained after Eq. (94). The last line in Eq. (107) symbolically represents the terms of the order of $O(\hat{X}_{x}^{2}$, $\hat{X}_{y}^{2}$, $\hat{X}_{XITM}^{2}$, $\hat{X}_{YITM}^{2}$, $\hat{D}_{v}\hat{X}_{x}$, $\hat{D}_{v}\hat{X}_{y}$, $\hat{D}_{v}\hat{X}_{XITM}$, $\hat{D}_{v}\hat{X}_{YITM}$, $\hat{A}\hat{X}_{x}$, $\hat{A}\hat{X}_{y}$, $\hat{A}\hat{X}_{XITM}$, $\hat{A}\hat{X}_{YITM})$. The input-output relation (107) is one of the main results of this paper.
To conduct the actual evaluation of the input-output relation (107), we have to evaluate ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ in some way. In the case of gravitational-wave detectors, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ are evaluated through the equations of motions for the end test masses as discussed in the next section.